OpenStudy (anonymous):

what is the weight of the contents of a right rectangular prism with length 15 ft, width 3ft, and height 3.5ft if the contents weigh 0.25 pounds per cubic ft.

jimthompson5910 (jim_thompson5910):

First find the volume: V = LWH = 15*3*3.5 = 157.5 So the volume of the prism is 157.5 cubic feet. Since the contents weigh 0.25 pounds per foot, this means that 1 cubic foot = 0.25 pounds, 2 cubic feet = 2*0.25 = 0.5 pounds, 3 cubic feet = 3*0.25 =0.75 pounds, etc So this means that 157.5 cubic feet weighs 157.5*0.25 = 39.375 pounds So the weight of the contents of the container is 39.375 pounds.